### v1

@@ 2.2 Identification
In Appendix A, we show that if the rule profile $\mathbf{f}$ satisfies Assumption 1, then the perceived dataset satisfies a cyclical consistency condition that generalizes the standard SARP (Strong Axiom of Revealed Preference) to our choice environment. This implication marks a fundamental difference with the standard revealed-preference approach. In classical utility-based models, cyclical consistency is imposed directly on the observed dataset $D$ and serves as a common weak denominator across broad families of theories, including expected utility theory and its generalizations. The data are required to look as if they were generated by a stable underlying ordering. In our framework, the raw dataset $D$ is not expected to satisfy such consistency with respect to the normative benchmark. Indeed, many canonical “anomalies” correspond precisely to monotonicity violations or to revealed-preference cycles under FSD. The discipline instead comes from a different place. Behavior is generated by menu-dependent rules that transform each problem into a perceived one, and the observed choice must be locally defensible as a strict FSD improvement in that perceived problem. Under FSD, this local discipline is strong enough to guarantee that the induced perceived dataset $\widetilde{D}(\mathbf{f})$ satisfies a strong cyclical consistency condition. The empirical content therefore shifts away from asking whether choices are globally consistent as observed, and toward asking whether they can be organized by a parsimonious collection of simple transformations that make them locally dominance-improving once perception is taken into account.
3 Examples
This section shows how familiar “anomalies” in risky choice can be viewed through our library. Each example illustrates which rules in $\mathcal{F}$ satisfy Local Discriminability.
3.1 FSD violation
Consider a single choice between lotteries $L^{1}$ and $L^{2}$ defined on three equally likely states $s\in\{1,2,3\}$. Lottery $L^{2}$ strictly first-order stochastically dominates $L^{1}$ for any $\varepsilon>0$: it yields a (weakly) higher payoff in every state and a strictly higher payoff in at least one state. Choosing $L^{1}$ therefore violates monotonicity with respect to $\geq_{\mathrm{FSD}}$.
Table 1: Violation of FSD and regret/salience rules
state $s$ 11 22 33
$p_{s}$ $1/3$ $1/3$ $1/3$
$L^{1}$ 2020 0 $10$
$L^{2}$ 0 $10+\varepsilon$ 2020
Empirically, such choices are not rare; see, e.g., bleichrodt2015. A standard explanation is regret. Under our regret rule REG, the severity of regret from choosing $L^{i}$ against $L^{j}$ is the largest state-wise payoff loss
$\rho^{i}(L^{1},L^{2})\ :=\ \max_{s\in S}\max\{\,z^{j}_{s}-z^{i}_{s},\,0\,\},$
where $(z^{1}_{s},z^{2}_{s})_{s\in S}$ denotes a common state-space representation of $(L^{1},L^{2})$ as in Section 2. In the menu $\{L^{1},L^{2}\}$, choosing $L^{2}$ exposes the decision maker to a potential forgone gain of $20$ in state $s=1$, whereas choosing $L^{1}$ exposes her only to a forgone gain of $10+\varepsilon$ in state $s=2$. Hence $\rho^{2}(L^{1},L^{2})>\rho^{1}(L^{1},L^{2})$ and REG strictly prefers $L^{1}$ to $L^{2}$.
The salience rule SAL delivers a distinct but related explanation. It computes the state-wise contrasts $c(z^{1}_{s},z^{2}_{s})$ and focuses attention on the most salient state. Here the largest payoff gap is again in state $s=1$ (2020 vs. 0), so SAL collapses the comparison to a sure $20$ under $L^{1}$ versus a sure 0 under $L^{2}$ and strictly prefers $L^{1}$. In terms of our library, this FSD violation can be explained by REG or SAL, while the other rules fail to yield a strict ranking (e.g., MMn yields 0 for both lotteries; MMx yields $20$ for both; DIS delivers the same disappointment severity for both; and MAP and PW do not discriminate).
3.2 Cyclical consistency (I)
Our next example, from bleichrodt2015 following loomes_sugden1987 and sugden1991, illustrates how behavior consistent with regret and salience can generate a revealed-preference cycle under $\geq_{\mathrm{FSD}}$ even though no FSD violation occurs at any single menu. There are four equiprobable states and four lotteries $L^{3},\dots,L^{6}$:
Table 2: Preference cycle (I)
state $s$ 11 22 33 44
$p_{s}$ $1/4$ $1/4$ $1/4$ $1/4$
$L^{3}$ 0 $10$ 2020 $30$
$L^{4}$ $30$ 0 $10$ 2020
$L^{5}$ 2020 $30$ 0 $10$
$L^{6}$ $10$ 2020 $30$ 0
The decision maker faces the menus
$\mathcal{A}=\big\{\{L^{3},L^{4}\},\ \{L^{4},L^{5}\},\ \{L^{5},L^{6}\},\ \{L^{6},L^{3}\}\big\},$
and we observe
$x_{1}=L^{4},\quad x_{2}=L^{5},\quad x_{3}=L^{6},\quad x_{4}=L^{3}.$
This is a classical example of cyclical consistency violation. [^7] With the notations of Appendix A, under $\geq_{\mathrm{FSD}}$, all four lotteries are pairwise incomparable: they share the same multiset of payoffs $\{0,10,20,30\}$ with equal probabilities. No single choice therefore violates FSD, and the strict revealed-preference relation $P$ (defined from $\geq_{\mathrm{FSD}}$) is empty. However, the weak relation $R$ from Section 2 satisfies $L^{4}\,R\,L^{3},\quad L^{5}\,R\,L^{4},\quad L^{6}\,R\,L^{5},\quad L^{3}\,R\,L^{6},$ so the objective dataset induces a nontrivial revealed-preference cycle (and hence fails the SARP-type condition in Definition 10, Appendix A). The observed choices are exactly reproduced by REG and SAL. Under REG, in each menu $\{L^{i},L^{j}\}$ the rejected lottery has regret severity $30$ while the chosen one has regret severity $10$. For instance, in $\{L^{3},L^{4}\}$ the regret from choosing $L^{3}$ is $30$ (state $1$), whereas the regret from choosing $L^{4}$ is $10$ (states 22-44). SAL behaves similarly: in each pair the most salient state is the one where one lottery yields $30$ and the other 0, and SAL chooses the $30$-paying lottery in that state, again generating the 4-cycle. By contrast, MMn, MMx, DIS, MAP, and PW do not strictly discriminate in these menus.
3.3 Common-consequence Allais paradox
We next connect the Allais common-consequence pattern to our rule library. Consider the standard pair of choice problems. In the first problem, subjects choose between
[…]
and in the second problem between
[…]
In this well-known example, many subjects choose $L^{7}$ over $L^{8}$ and $L^{10}$ over $L^{9}$, violating the independence axiom of expected utility. Within our rule library, SAL and REG provide natural explanations. A salience thinker focuses on the possibility of winning \$2,500 versus \$0 in the first problem, and on the possibility of receiving \$0 versus \$2,400 in the second. A regret-based thinker compares worst forgone outcomes across the two menus. In particular, under REG, in the first choice problem the decision maker prefers $L^{7}$ because choosing $L^{8}$ entails a worst forgone gain of \$2,500 (relative to the state in which $L^{7}$ pays \$2,500 and $L^{8}$ pays \$0). In the second choice problem, the sure \$2,400 outcome has strictly lower regret severity.
3.4 Common-ratio Allais paradox
Consider now the common-ratio pair. In the first problem, subjects choose between
[…]
and in the second problem between
[…]
The classical pattern is $L^{11}$ chosen over $L^{12}$ and $L^{13}$ chosen over $L^{14}$. No single rule from $\mathcal{F}$ can explain both menus while satisfying Local Discriminability. PW, for instance, is often invoked for common-ratio patterns, but here its coarse coding of probabilities does not yield a strict ranking in the first problem. The modal-payoff rule MAP, on the other hand, does. In $\{L^{11},L^{12}\}$ the modal payoff for $L^{11}$ is \$3,000 (probability $0.95$), while the modal payoff for $L^{12}$ is \$0 (probability $0.525$), so MAP strictly prefers $L^{11}$. In the second menu, SAL and REG both favor $L^{13}$: they focus on the contrast between \$6,000 and \$0, or on the worst forgone loss from not choosing $L^{13}$. Accordingly, a locally discriminating rationalization uses MAP in the first problem and SAL or REG in the second.
3.5 The reflection effect
We now turn to the reflection effect: risk aversion over fair gains and risk seeking over fair losses. Consider two 50-50 choices over gains and losses. In the gain domain:
$L^{15}=\begin{cases}100&\text{with prob. }1,\end{cases}\qquad L^{16}=\begin{cases}200&\text{with prob. }0.5,\\ 0&\text{with prob. }0.5,\end{cases}$
and in the loss domain:
$L^{17}=\begin{cases}-100&\text{with prob. }1,\end{cases}\qquad L^{18}=\begin{cases}0&\text{with prob. }0.5,\\ -200&\text{with prob. }0.5.\end{cases}$
All gambles are fair: $\mathbb{E}[L^{15}]=\mathbb{E}[L^{16}]=100$ and $\mathbb{E}[L^{17}]=\mathbb{E}[L^{18}]=-100$. A common pattern selects $L^{15}$ and $L^{18}$.
In our library, REG predicts indifference in both problems: regret severity is the same for the two lotteries in each menu, so REG cannot generate the reflection effect. SAL, by contrast, is consistent with the observed pattern. In $\{L^{15},L^{16}\}$, the largest contrast is between receiving \$100 under $L^{15}$ and \$0 under $L^{16}$, leading to risk aversion. In $\{L^{17},L^{18}\}$, the most salient gap is between a loss of \$100 and no loss (0), pushing choice toward the risky $L^{18}$. The key driver is diminishing sensitivity in the contrast function: the gap between 0 and $\pm 100$ has higher contrast than that between $\pm 100$ and $\pm 200$.
3.6 Cyclical consistency (II)
Our next example shows a pure salience-driven violation of strong cyclical consistency. Consider three equiprobable states and three lotteries $L^{19},L^{20},L^{21}$:
Table 3: Three-lottery salience cycle
state $s$ 11 22 33
$p_{s}$ $1/3$ $1/3$ $1/3$
$L^{19}$ 0 0 2020
$L^{20}$ $10$ 88 0
$L^{21}$ 0 $30$ 0
The menus are
$\mathcal{A}=\big\{\{L^{19},L^{20}\},\ \{L^{20},L^{21}\},\ \{L^{21},L^{19}\}\big\},$
and suppose the choices are $L^{19}$, $L^{20}$, and $L^{21}$ in respectively the first, second, and third choice problem. [^8] Since $L^{19}\neq L^{21}$, the dataset violates the strong cyclical consistency condition of Appendix A. Under SAL, in $\{L^{19},L^{20}\}$ the most salient gap is in $s=3$ (2020 vs. 0), so $L^{19}$ is chosen. In $\{L^{20},L^{21}\}$, the contrast in $s=1$ is $c(10,0)=10/11$, larger than the contrast in $s=2$, $c(8,30)=22/39$, so $L^{20}$ is chosen. In $\{L^{21},L^{19}\}$, the largest gap is in $s=2$ ($30$ vs. 0), so $L^{21}$ is chosen. REG does not reproduce the same cycle because regret severity for $L^{20}$ in $\{L^{20},L^{21}\}$ comes from state $2$ (foregone $30$), which exceeds the regret severity for $L^{21}$; REG therefore prefers $L^{21}$ to $L^{20}$ in that menu. MMn, MMx, and DIS also treat the three lotteries more symmetrically and do not yield a 3-cycle. This example isolates a pattern that is salience-based.
3.7 Mixture reversal
Finally, we consider a mixture reversal: a preference reversal induced by mixing both options with a common baseline. In the first problem, the decision maker chooses between a safe and a risky gain:
$L^{24}=\begin{cases}200&\text{with prob. }1\end{cases}\qquad\text{vs}\qquad L^{25}=\begin{cases}300&\text{with prob. }0.8\\ 0&\text{with prob. }0.2\;.\end{cases}$
In the second problem, both lotteries are mixed with a sure-zero baseline $L^{0}=(0\ \text{with prob. }1)$ with weight $\lambda=0.6$, $L^{26}=\lambda L^{24}+(1-\lambda)L^{0}$ and $L^{27}=\lambda L^{25}+(1-\lambda)L^{0}$:
[…]
Suppose that we observe the choices $L^{24}$ and $L^{27}$, the standard mixture reversal pattern.
Within our library, DIS gives a natural explanation. In the first problem, $L^{24}$ has modal payoff $ref(L^{24})=200$ and no lower outcome, so $D(L^{24})=0$. Lottery $L^{25}$ has modal payoff $ref(L^{25})=300$ and a lower outcome 0, so $D(L^{25})=c(300,0)>0$, and DIS prefers the sure outcome. In the second problem, $L^{26}$ has modal outcome $200$ and lower outcome 0, so $D(L^{26})=c(200,0)>0$, whereas $L^{27}$ has modal outcome 0 (probability $0.52$) and no strictly lower outcome, so $D(L^{27})=0$. DIS therefore reverses its ranking and now prefers $L^{27}$. The mixture with $L^{0}$ flips the modal (reference) payoff for the risky option but not for the safe one, eliminating disappointment for $L^{27}$ and introducing it for $L^{26}$.
SAL and REG can also rationalize the reversal. In the first problem, both focus on the event where $L^{25}$ yields 0 and $L^{24}$ yields $200$, favoring the safe option; in the second, they focus on the event where $L^{27}$ yields $300$ while $L^{26}$ yields 0, favoring the risky option. By contrast, MMn selects the safe option in the first problem but is indifferent in the second (both minima are zero), MMx always prefers the risky option, and MAP predicts $L^{25}$ in the first problem and $L^{26}$ in the second.
Discussion
Each example isolates which rules in the library $\mathcal{F}$ can explain the observed choices while satisfying Local Discriminability. Some patterns are compatible with several distinct rules (e.g., the common-consequence pair can be explained by either salience or regret), while others are not. For instance, the reflection effect separates salience from regret, since REG predicts indifference whereas SAL delivers the observed switch from risk aversion over gains to risk seeking over losses; the common-ratio pair forces the use of the modal-payoff rule MAP in the first menu and rules such as SAL or REG in the second; and the three-lottery cycle (Cyclical consistency II) is essentially salience-specific. Whenever such patterns are present in the data, they mechanically restrict which rules can be used at the corresponding menus in any admissible profile.
Taken together, the examples show that a small, nonparametric library of rules can organize a broad range of well-known behaviors in risky choice, and that different patterns load on different subsets of rules. In the next section, we move from these qualitative illustrations to a quantitative complexity measure that asks how parsimoniously a whole dataset explained by such rules, and how much of this parsimony is carried by specific rules.
@@ 4.1 Admissible profiles and rule assignments
Let $\mathcal{C}(D;\geq_{FSD},\mathcal{F})$ be the set of rule assignments $\mathbf{y}=\{y_{t,f}\}$ defined on $\mathcal{T}$ such that, for each $t\in\mathcal{T}$, exactly one rule is used and that rule is strictly discriminating at $t$:
@@
Elements of $\mathcal{C}(D;\geq_{FSD},\mathcal{F})$ are called admissible rule assignments.
@@ 4.2 The Maximum Rule Concentration Index
$\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})\ :=\ \max_{\mathbf{y}\in\mathcal{C}(D;\geq_{FSD},\mathcal{F})}\mathrm{HHI}(\mathbf{y}),$
and the effective number of rules is $N_{\mathrm{eff}}:=1/\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})$.
@@ 2. $i$
    If for some $t\in\mathcal{T}$ there is a unique strictly discriminating rule $f$ (i.e., $\mathcal{F}_{t}^{\mathrm{strict}}=\{f\}$), then every maximizer $\mathbf{y}^{\star}$ of $\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})$ satisfies $y^{\star}_{t,f}=1$.
@@
Because admissible assignments and MRCI–maximizers need not be unique, we look at how much concentration can be preserved when rules are deleted. Fix a finite family $\mathcal{K}$ of permutations of $\mathcal{F}$. For each order $k\in\mathcal{K}$, we construct a Deletion–Stable Set $E^{(k)}$ by iteratively removing rules $f^{(k)}_{i}$ if the MRCI on the restricted library remains equal to the original optimum $H^{\star}=\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})$. The final set $E^{(k)}$ is a minimal library that maintains the coherence of the originally explained behavior.
@@
$\phi(f)\ :=\ \frac{\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})\;-\;\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F}\setminus\{f\})}{\operatorname{MRCI}(D;\geq_{FSD},\mathcal{F})}\in[0,1].$
@@
4.3 Examples, revisited
To illustrate these notions, it is useful to revisit the examples from Section 3. For each pattern we can identify (i) which rules are locally admissible at each menu, (ii) whether a one-rule explanation exists, and (iii) which rules become indispensable once we pool examples.
Single-example datasets.
For most canonical patterns taken in isolation (e.g., the FSD violation, the two cyclical examples, and the common-consequence pair), the admissible set $\mathcal{C}(D)$ contains assignments that use only one non-attention rule from $\mathcal{F}$ across all observations. For instance, in the FSD-violation example, either REG or SAL can be used at the single menu in a way that satisfies Local Discriminability. For such small datasets, MRCI therefore cannot distinguish sharply between candidate rules: several heuristics are perfect substitutes for explaining the pattern under Local Discriminability, which leads to $\phi(f)=0$ for each of them and stability scores that spread symmetrically across the admissible rules. Isolated anomalies mainly tell us that some simple rule can account for the behavior while preserving perceived rationality; they do not yet pin down which rule is responsible.
A pooled illustration.
The value of MRCI and the rule-usage diagnostics becomes more apparent when we pool multiple examples, as in standard datasets. Let $D_{\mathrm{pool}}$ collect the sixteen observations spanning: the FSD violation; the four-lottery cycle (Cyclical consistency I); the common-consequence Allais pair; the common-ratio pair; the reflection effect (two menus); the three-lottery salience cycle (Cyclical consistency II); and the mixture reversal (two menus). For each observation, Table 4 summarizes which non-attention rules in $\mathcal{F}$ are locally admissible (i.e., belong to $\mathcal{F}_{t}^{\mathrm{strict}}$) once Local Discriminability is imposed.
Table 4: Summary: examples and locally admissible non-attention rules
Example Admissible rules in $\mathcal{F}_{t}^{\mathrm{strict}}\setminus\{\textsc{A1},\textsc{A2}\}$
FSD violation REG, SAL
Cyclical consistency (I) REG, SAL
Common-consequence REG, SAL
Common-ratio $\{L^{11},L^{12}\}$: MAP; $\{L^{13},L^{14}\}$: REG, SAL
Reflection effect SAL
Cyclical consistency (II) SAL
Mixture reversal DIS, REG, SAL
In this pooled dataset $D_{\mathrm{pool}}$, SAL is locally admissible in fifteen out of sixteen observations; the only exception is the first common-ratio menu $\{L^{11},L^{12}\}$, where only MAP and one attention rule are locally discriminating. There therefore exists an admissible assignment that uses SAL at all observations where it is admissible and MAP at the remaining one. Thus every maximizer $\mathbf{y}^{\star}$ of $\mathrm{HHI}$ satisfies

### v2

@@ 2.2 Identification
In Appendix B, we show that if the rule profile $\mathbf{f}$ satisfies Assumption 1, then the perceived dataset satisfies a cyclical consistency condition that generalizes the standard SARP (Strong Axiom of Revealed Preference) to our choice environment. This implication marks a fundamental difference with the standard revealed-preference approach. In classical utility-based models, cyclical consistency is imposed directly on the observed dataset $D$ and serves as a common weak denominator across broad families of theories, including expected utility theory and its generalizations. The data are required to look as if they were generated by a stable underlying ordering. In our framework, the raw dataset $D$ is not expected to satisfy such consistency with respect to the normative benchmark. Indeed, many canonical “anomalies” correspond precisely to monotonicity violations or to revealed-preference cycles under FSD. The discipline instead comes from a different place. Behavior is generated by menu-dependent rules that transform each problem into a perceived one, and the observed choice must be locally defensible as a strict FSD improvement in that perceived problem. Under FSD, this local discipline is strong enough to guarantee that the induced perceived dataset $\widetilde{D}(\mathbf{f})$ satisfies a strong cyclical consistency condition. The empirical content therefore shifts away from asking whether choices are globally consistent as observed, and toward asking whether they can be organized by a parsimonious collection of simple transformations that make them locally dominance-improving once perception is taken into account.
@@ 3.1 Admissible profiles and rule assignments
Let $\mathcal{C}(D;\geq_{\mathrm{FSD}},\mathcal{F})$ be the set of rule assignments $\mathbf{y}=\{y_{t,f}\}$ defined on $\mathcal{T}$ such that, for each $t\in\mathcal{T}$, exactly one rule is used and that rule is strictly discriminating at $t$:
@@
Elements of $\mathcal{C}(D;\geq_{\mathrm{FSD}},\mathcal{F})$ are called admissible rule assignments.
@@ 3.2 The Maximum Rule Concentration Index
$\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})\ :=\ \max_{\mathbf{y}\in\mathcal{C}(D;\geq_{\mathrm{FSD}},\mathcal{F})}\mathrm{HHI}(\mathbf{y}),$
and the effective number of rules is $N_{\mathrm{eff}}:=1/\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})$.
@@ 2. $i$
    If for some $t\in\mathcal{T}$ there is a unique strictly discriminating rule $f$ (i.e., $\mathcal{F}_{t}^{\mathrm{strict}}=\{f\}$), then every maximizer $\mathbf{y}^{\star}$ of $\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})$ satisfies $y^{\star}_{t,f}=1$.
@@
Because admissible assignments and MRCI–maximizers need not be unique, we look at how much concentration can be preserved when rules are deleted. Fix a finite family $\mathcal{K}$ of permutations of $\mathcal{F}$. For each order $k\in\mathcal{K}$, we construct a Deletion–Stable Set $E^{(k)}$ by iteratively removing rules $f^{(k)}_{i}$ if the MRCI on the restricted library remains equal to the original optimum $H^{\star}=\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})$. The final set $E^{(k)}$ is a minimal library that maintains the coherence of the originally explained behavior.
@@
$\phi(f)\ :=\ \frac{\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})\;-\;\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F}\setminus\{f\})}{\operatorname{MRCI}(D;\geq_{\mathrm{FSD}},\mathcal{F})}\in[0,1].$
@@
3.3 Examples
We keep this section deliberately short. We select four canonical patterns, and we compute the $\operatorname{MRCI}$ by hand on a toy dataset obtained by pooling these examples. The point is not to provide a taxonomy of anomalies, but to illustrate (i) how the local dominance screen works menu-by-menu, and (ii) how the $\operatorname{MRCI}$ aggregates those local admissibility sets into a global, maximally concentrated explanation.
Example 1: An FSD violation
Consider a single choice between lotteries $L^{1}$ and $L^{2}$ defined on three equally likely states $s\in\{1,2,3\}$:
Table 1: Violation of FSD
state $s$ 11 22 33
$p_{s}$ $1/3$ $1/3$ $1/3$
$L^{1}$ 2020 0 $10$
$L^{2}$ 0 $10+\varepsilon$ 2020
Lottery $L^{2}$ strictly first-order stochastically dominates $L^{1}$ for any $\varepsilon>0$, so choosing $L^{1}$ violates monotonicity under $\geq_{\mathrm{FSD}}$. Under REG, however, choosing $L^{2}$ entails a larger worst foregone gain (state $1$) than choosing $L^{1}$ (state $2$), so REG strictly prefers $L^{1}$. Under SAL, the most salient state is again $s=1$ (largest contrast), so SAL also strictly prefers $L^{1}$. Hence, at this menu,
$\mathcal{F}^{\mathrm{strict}}_{1}\setminus\{\textsc{A1},\textsc{A2}\}=\{\textsc{REG},\textsc{SAL}\}.$
Example 2: A common-ratio Allais Paradox
Consider the common-ratio problems:
$L^{3}=\begin{cases}3{,}000&\text{w.p. }0.9\\ 0&\text{w.p. }0.1\end{cases}\quad\text{vs}\quad L^{4}=\begin{cases}6{,}000&\text{w.p. }0.45\\ 0&\text{w.p. }0.55\end{cases}$
and
$L^{5}=\begin{cases}6{,}000&\text{w.p. }0.001\\ 0&\text{w.p. }0.999\end{cases}\quad\text{vs}\quad L^{6}=\begin{cases}3{,}000&\text{w.p. }0.002\\ 0&\text{w.p. }0.998.\end{cases}$
A classical common-ratio pattern selects $L^{3}$ over $L^{4}$ and $L^{5}$ over $L^{6}$. Under our coarse library, the first menu is locally rationalized by MAP: the modal payoff of $L^{3}$ is 3,000 (probability $0.9$) while the modal payoff of $L^{4}$ is 0 (probability $0.55$), so MAP strictly prefers $L^{3}$. In the second menu, SAL or REG can justify choosing $L^{5}$ (focusing on the salient 6,000 vs. 0 contrast or on worst foregone losses). Hence,
$\mathcal{F}^{\mathrm{strict}}_{2}\setminus\{\textsc{A1},\textsc{A2}\}=\{\textsc{MAP}\},\qquad\mathcal{F}^{\mathrm{strict}}_{3}\setminus\{\textsc{A1},\textsc{A2}\}=\{\textsc{REG},\textsc{SAL}\}.$
Example 3: A mixture reversal
Consider a mixture reversal with baseline $L^{0}=(0\ \text{w.p. }1)$. In the first menu,
$L^{7}=\begin{cases}200&\text{w.p. }1\end{cases}\quad\text{vs}\quad L^{8}=\begin{cases}300&\text{w.p. }0.8\\ 0&\text{w.p. }0.2,\end{cases}$
and in the second menu both lotteries are mixed with $L^{0}$ using weight $\lambda=0.6$, $L^{9}=\lambda L^{7}+(1-\lambda)L^{0}$ and $L^{10}=\lambda L^{8}+(1-\lambda)L^{0}$:
[…]
The empirical pattern is that some subjects choose the safe option $L^{7}$ in the first menu but reverse after mixing, choosing $L^{10}$. In our library, this pair can be locally rationalized by SAL, REG, and DIS (contrast-, regret-, or disappointment- driven comparisons). Accordingly,
$\mathcal{F}^{\mathrm{strict}}_{4}\setminus\{\textsc{A1},\textsc{A2}\}=\mathcal{F}^{\mathrm{strict}}_{5}\setminus\{\textsc{A1},\textsc{A2}\}=\{\textsc{DIS},\textsc{REG},\textsc{SAL}\}.$
Example 4: A reflection effect
Consider two 50-50 choices over gains and losses:
$L^{11}=\begin{cases}100&\text{with prob. }1,\end{cases}\qquad L^{12}=\begin{cases}200&\text{with prob. }0.5,\\ 0&\text{with prob. }0.5,\end{cases}$
and in the loss domain:
$L^{13}=\begin{cases}-100&\text{with prob. }1,\end{cases}\qquad L^{14}=\begin{cases}0&\text{with prob. }0.5,\\ -200&\text{with prob. }0.5.\end{cases}$
The empirical patter is that some subjects choose $L^{12}$ and $L^{14}$. While there are several combinations of different rules that explain the observed patterns, only SAL is consistent with both observations. In $\{L^{11},L^{12}\}$, the largest contrast is between receiving 100 under $L^{11}$ and 0 under $L^{12}$, leading to risk aversion. In $\{L^{13},L^{14}\}$, the most salient gap is between a loss of 100 and no loss (0), pushing choice toward the risky $L^{14}$. [^7] REG predicts indifference in both problems, as the regret severity is the same for the two lotteries in each menu. Accordingly,
[…]
$\mathcal{F}^{\mathrm{strict}}_{7}\setminus\{\textsc{A1},\textsc{A2}\}=\{\textsc{SAL},\textsc{MMx}\}.$
Computing the $\operatorname{MRCI}$ on the pooled three-example dataset.
Let $D_{\mathrm{toy}}$ pool these three examples into $T=5$ observations: one FSD-violation menu ($t=1$), two common-ratio menus ($t=2,3$), and two mixture-reversal menus ($t=4,5$). The strictly admissible non-attention rules at each observation are:
$\mathcal{F}^{\mathrm{strict}}_{1}\setminus\{\textsc{A1},\textsc{A2}\}$ $=\{\textsc{REG},\textsc{SAL}\},$
$\mathcal{F}^{\mathrm{strict}}_{2}\setminus\{\textsc{A1},\textsc{A2}\}$ $=\{\textsc{MAP}\}$
$\mathcal{F}^{\mathrm{strict}}_{3}\setminus\{\textsc{A1},\textsc{A2}\}$ $=\{\textsc{REG},\textsc{SAL}\}$
$\mathcal{F}^{\mathrm{strict}}_{4}\setminus\{\textsc{A1},\textsc{A2}\}$ $=\mathcal{F}^{\mathrm{strict}}_{5}\setminus\{\textsc{A1},\textsc{A2}\}=\{\textsc{DIS},\textsc{REG},\textsc{SAL}\}$
$\mathcal{F}^{\mathrm{strict}}_{6}\setminus\{\textsc{A1},\textsc{A2}\}$ $=\{\textsc{SAL},\textsc{DIS},\textsc{MMn}\}$
$\mathcal{F}^{\mathrm{strict}}_{7}\setminus\{\textsc{A1},\textsc{A2}\}$ $=\{\textsc{SAL},\textsc{MMx}\}.$
An admissible explanation assigns to each observation one rule $f_{t}\in\mathcal{F}^{\mathrm{strict}}_{t}$ (and recall A1/A2 are always available). To maximize concentration, we seek an assignment that uses a single rule as often as possible. Here, SAL is admissible in $6$ out of $7$ observations (all except the first common-ratio menu $t=2$), so there exists an admissible assignment that uses SAL six times and uses MAP once. The resulting shares are $s_{\textsc{SAL}}=6/7$ and $s_{\textsc{MAP}}=1/7$ (all other non-attention rules have share 0 under this assignment), and the induced concentration is
$\operatorname{HHI}=\sum_{f\in\mathcal{F}}s_{f}^{2}=\Big(\tfrac{6}{7}\Big)^{2}+\Big(\tfrac{1}{7}\Big)^{2}\approx 0.76.$
This value is maximal. The reason is that menu $t=2$ admits only MAP among non-attention rules, so any admissible explanation must assign a non-SAL rule at least once. Since SAL is admissible at all other menus, the best possible concentration is achieved by assigning SAL everywhere it is admissible and MAP at $t=2$. Therefore,
$\operatorname{MRCI}(D_{\mathrm{toy}};\geq_{\mathrm{FSD}},\mathcal{F})\approx 0.76.$
This hand calculation illustrates the logic of $\operatorname{MRCI}$ in a nutshell. Local admissibility sets overlap substantially. Several menus admit SAL and REG, but the $\operatorname{MRCI}$ resolves this overlap by choosing, menu-by-menu, the assignment that yields the most concentrated global organization. The toy calculation also makes the logic of our diagnostics transparent once we remember that attention rules are always available. In particular, if we delete MAP from the library, the forced common-ratio menu $t=2$ can be rationalized by an attention default (A1 or A2), while the other menus can still be assigned to SAL. Hence the maximal concentration is unchanged, so the Concentration Gain of MAP is 0. By contrast, if we delete SAL, we lose the only rule that (in this toy dataset) is simultaneously admissible at all of $t\in\{1,3,4,5,6,7\}$. The best we can then do is to concentrate on one of the remaining rules that has the widest overlap, REG, which is admissible at $t\in\{1,3,4,5\}$ but competes with attention defaults at those menus. Here, if no single attention rule is feasible at $t\in\{1,3,4,5\}$, a conservative assignment leads to